# ContextMRI: Enhancing Compressed Sensing MRI through Metadata Conditioning

### v1

@@ Discussion
First, note that CLIP [38], the text encoder used for ContextMRI, was used without fine-tuning it to MRI image-text pairs. Naturally, it may not understand the texts, potentially making it harder for the diffusion model to learn the relation between the given metadata and images. An interesting future study will involve designing and training a text encoder specific to MRI or incorporating large language models (LLM) that were pre-trained on medical data as a text encoder.
Second, a key limitation of leveraging text-to-image diffusion models for CS-MRI lies in the heterogeneity of metadata across datasets and clinical settings. Available metadata often varies depending on the source of the data. For instance, datasets such as SKM-TEA [35] provide patient-specific details like age and sex, whereas fastMRI [39] primarily includes imaging parameters and lacks patient demographic information. This variability complicates the design of a unified structural format for input metadata, which is crucial for effectively conditioning the generative model.
@@ 0.6 Data curation
Our main dataset collection is based on the fastMRI [39] data on the knee and brain. From the raw $k$-space data of varying size, we compute the inverse Fourier transform, and center-crop all images to resolution $320\times 320$, so that the unwanted boundaries are eliminated. We take the Fourier transform back to the $k$-space and estimate the sensitivity maps through ESPiRiT [29] to compute the MVUE estimate. We normalize the images by dividing the complex image with 99% quantile, which approximately provides the range of $[-1.5,1.5]$ for both the real and the imaginary channels. Our training set consists of 167,375 slices across 8,344 volumes.
For the patient demographic experiment, we fine-tune our model using the SKM-TEA [35] dataset, where the images are of $512\times 512$ resolution. We use the same data preprocessing steps as in the fastMRI case and directly use the sensitivity maps provided in the dataset. The training consists of 25,731 slices of images across 139 volumes.
@@ 0.7 Metadata conditioning
For the metadata, we consider the following entries: anatomy, slice, contrast, pathology, sequence, TR, TE, TI, flip angle. These are structured into a string, similar to the following example: “Knee, Slice 19, PDFS, Pathology: Displaced Meniscal Tissue, Meniscus Tear, Bone-Subchondral edema, TR: 3150, TE: 33, TI: 100, Flip angle: 150”. All the metadata except for the pathology information is extracted from the metadata provided directly in the fastMRI database. See Tab. 2, Fig. 1, Fig. 2, and Fig. 3 for the metadata conditions used, and their statistics. In the fastMRI+ [40] database, there are annotations with bounding boxes whenever a pathology can be seen in the specified slice. We ignore the location of the bounding box and only append the existing types of pathologies that can be found in the slice, separated with a comma. When no pathologies are found, we do not append this trait. To enhance the robustness for the cases where some of the metadata might not be provided during inference, we randomly drop the MR imaging parameters TR, TE, TI, and Flip angle with 50% probability during training.
@@ 0.8 Diffusion model training
Consider a continuous sequence of densities $p_{t}(\bm{x})$, where $t$ indicates time in a diffusion process. Let $\bm{x}_{0}\sim p_{0}(\bm{x})$ be the data distribution, and $\bm{x}_{T}=\mathcal{N}(0,I_{d})$ be the reference Gaussian normal distribution. In diffusion models [2, 1], the data distribution is gradually corrupted with Gaussian noise with the perturbation kernel $p(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{0};0,t^{2}I_{d})$, until it approximates a standard normal distribution. Diffusion models learn the denoising generative reverse process, i.e. creating data from noise. This process is governed by the score function $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$ [41], which is estimated through denoising score matching $\bm{s}_{\theta}(\bm{x}_{t})\approx\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$ [41]. Text conditional diffusion models [10, 42] are trained with image-text pairs with random dropping so that it is possible to both sample from $p(\bm{x})$ by using $\bm{\epsilon}_{\theta}(\bm{x}_{t})$ or from $p(\bm{x}|{\bm{c}})$ by using $\bm{\epsilon}_{\theta}(\bm{x}_{t},{\bm{c}})$, where ${\bm{c}}$ is the text embedding. Further, to emphasize the conditioning signal, another widely used technique is classifier free guidance (CFG) [13], where $\bm{\epsilon}_{\theta}^{\gamma}(\bm{x}_{t},{\bm{c}}):=\bm{\epsilon}_{\theta}(\bm{x}_{t})+\gamma(\bm{\epsilon}_{\theta}(\bm{x}_{t},{\bm{c}})-\bm{\epsilon}_{\theta}(\bm{x}_{t}))$ is used to sample from the sharpened posterior $p(\bm{x})p({\bm{c}}|\bm{x})^{\gamma}$.
As the loss function, we use the epsilon-matching loss function commonly used in DDPMs [2] and employ a fixed variance setting rather than learning the variance. The model parameters are optimized for 50 epochs using the AdamW optimizer with a learning rate of 0.0001. All experiments were conducted using PyTorch 2.1 in Python with CUDA 12.1, running on 8 NVIDIA H100 GPUs (80GB). The effective batch size was set to 64.
@@ 0.9 Diffusion model architecture
We utilize a pixel-level diffusion model for both unconditional and metadata-conditional generation. The decision to employ pixel-level diffusion, rather than latent-space diffusion via a Variational Autoencoder (VAE) [43], stems from the inherent limitations of VAEs. Specifically, VAE-based reconstruction often compromises high-frequency details [44], and solving inverse problems in the latent space requires decoding to ensure data consistency. While such methods [15, 45, 16] are theoretically feasible, they frequently fail in practice and lack consensus as a reliable solution. Consequently, we opted to perform generation directly in pixel space.
To train the diffusion model on MRI complex values, we separate the real and imaginary parts, concatenating them along the channel dimension to create a 2-channel input. This results in a 2D input of size 2 × 320 × 320, which is used during training. Our network is adapted from the DeepFloyd IF [42] pixel-level text-to-image diffusion model, with modifications to accommodate our input format by adjusting the number of channels in the layers. Unlike natural image generation, the textual metadata associated with MRI data is significantly smaller. Therefore, we use the CLIP text encoder instead of the T5-XXL encoder employed in DeepFloyd IF, offering a more compact and memory-efficient representation of the metadata. During training, the CLIP text encoder is frozen, and only the parameters of the UNet modules are learnable.
@@ 0.10 Inverse problem solving with diffusion models
where for the multi-coil MRI case [46], the forward model is defined in parallel for each measurement coil
@@
to write the score function of the posterior in terms of the prior score function. A vast literature on DIS can be structured as different ways of approximating the intractable time-dependent likelihood $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ [47]. A few recent works [15, 16, 17] tried to incorporate text conditions through Stable Diffusion [10]. Among these, P2L [15] is prohibitively slow and is highly unstable when used with CFG. TReg [16] is tailored for severe degradations and produces cartoon-like artifacts. Dreamsampler [17] suggests tailored algorithms for specific types of inverse problems. Generally, setting moderately high values of CFG leads to a steep decrease in the stability of all the algorithms. None of the existing methods have shown the stability and performance required to extend the method to the medical imaging setting, where this is crucial.
We opt for a simple, established baseline for our choice of diffusion model-based inverse problem solver. For this, we choose DDS [9], which is known to be especially robust to large-scale medical imaging inverse problems. Specifically, with a variance preserving form, DDS with DDIM sampling [48] can be described as iterating the following steps for $t=T,T-1,\dots,1$ initialized with random Gaussian noise.
^$\hat{}\bm{x}_{0|t}$ $=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}^{\gamma}(\bm{x}_{t},{\bm{c}})\right)$ (4)
^$\hat{}\bm{x}^{\prime}_{0|t}$ $=\operatorname*{arg\,min}_{\bm{x}}\frac{\xi}{2}\|\bm{y}-A\bm{x}\|_{2}^{2}+\frac{1}{2}\|\bm{x}-\hat{}\bm{x}_{0|t}\|_{2}^{2}$ (5)
@@
where $\sigma_{t}:=\eta\sqrt{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ with a constant hyperparameter $\eta=0.8$. From a high-level viewpoint, skipping the data consistency optimization step in Eq. 5 would lead to DDIM sampling from the prior distribution $p(\bm{x})$. In order to sample from the approximate posterior, DDS modulates the posterior mean ^$\hat{}\bm{x}_{0|t}={{\mathbb{E}}}[\bm{x}_{0}|\bm{x}_{t}]$ with data consistency steps to approximate the conditional posterior mean ^$\hat{}\bm{x}^{\prime}_{0|t}\approx{{\mathbb{E}}}[\bm{x}_{0}|\bm{x}_{t},\bm{y}]$. Specifically, for the optimization problem in Eq. 5, the key of DDS is to use few-step conjugate gradient steps. In our case, we use 5-step update with $\xi=5.0$.
Our method is tested on two different undersampling patterns - uniform1D as implemented in fastMRI [39], and poisson2D as implemented in sigpy [49]. For uniform1D, we test two different acceleration factors: ×$\times 4$ with 8% autocalibrating signal (ACS) region in the center, and ×$\times 8$ with 4% ACS region. For poisson2D, we test ×$\times 8$ and ×$\times 15$ acceleration. For quantitative evaluation, we use three standard metrics - PSNR, SSIM, and LPIPS.
Data Availability
All data to reproduce ContextMRI are open-sourced in the fastMRI [39], fastMRI+ [40], and SKM-TEA [35] database.
Code Availability
The source code for ContextMRI, along with the code for preprocessing the metadata, will be open-sourced upon publication
Supplementary Information
Supplementary Figure 1: Metadata statistics of fastMRI knee train set
Supplementary Figure 2: Metadata statistics of fastMRI brain train set
Supplementary Figure 3: Metadata statistics of SKM-TEA dataset

### v2

@@ 3 Discussion
First, note that CLIP [radford2021learning], the text encoder used for ContextMRI, was used without fine-tuning it to MRI image-text pairs. Naturally, it may not understand the texts, potentially making it harder for the diffusion model to learn the relation between the given metadata and images. An interesting future study will involve designing and training a text encoder specific to MRI or incorporating large language models (LLM) that were pre-trained on medical data as a text encoder.
Second, a key limitation of leveraging text-to-image diffusion models for CS-MRI lies in the heterogeneity of metadata across datasets and clinical settings. Available metadata often varies depending on the source of the data. For instance, datasets such as SKM-TEA [desai2022skm] provide patient-specific details like age and sex, whereas fastMRI [zbontar2018fastmri] primarily includes imaging parameters and lacks patient demographic information. This variability complicates the design of a unified structural format for input metadata, which is crucial for effectively conditioning the generative model.
@@ 4.1 Data curation
Our main dataset collection is based on the fastMRI [zbontar2018fastmri] data on the knee and brain. From the raw $k$-space data of varying size, we compute the inverse Fourier transform, and center-crop all images to resolution $320\times 320$, so that the unwanted boundaries are eliminated. We take the Fourier transform back to the $k$-space and estimate the sensitivity maps through ESPiRiT [uecker2014espirit] to compute the MVUE estimate. We normalize the images by dividing the complex image with 99% quantile, which approximately provides the range of $[-1.5,1.5]$ for both the real and the imaginary channels. Our training set consists of 167,375 slices across 8,344 volumes.
For the patient demographic experiment, we fine-tune our model using the SKM-TEA [desai2022skm] dataset, where the images are of $512\times 512$ resolution. We use the same data preprocessing steps as in the fastMRI case and directly use the sensitivity maps provided in the dataset. The training consists of 25,731 slices of images across 139 volumes.
@@ 4.2 Metadata conditioning
For the metadata, we consider the following entries: anatomy, slice, contrast, pathology, sequence, TR, TE, TI, flip angle. These are structured into a string, similar to the following example: “Knee, Slice 19, PDFS, Pathology: Displaced Meniscal Tissue, Meniscus Tear, Bone-Subchondral edema, TR: 3150, TE: 33, TI: 100, Flip angle: 150”. All the metadata except for the pathology information is extracted from the metadata provided directly in the fastMRI database. See Tab. 2, Fig. 6, Fig. 7, and Fig. 8 for the metadata conditions used, and their statistics. In the fastMRI+ [zhao2021fastmri+] database, there are annotations with bounding boxes whenever a pathology can be seen in the specified slice. We ignore the location of the bounding box and only append the existing types of pathologies that can be found in the slice, separated with a comma. When no pathologies are found, we do not append this trait. To enhance the robustness for the cases where some of the metadata might not be provided during inference, we randomly drop the MR imaging parameters TR, TE, TI, and Flip angle with 50% probability during training.
@@ 4.3 Diffusion model training
Consider a continuous sequence of densities $p_{t}({\bm{x}})$, where $t$ indicates time in a diffusion process. Let ${\bm{x}}_{0}\sim p_{0}({\bm{x}})$ be the data distribution, and ${\bm{x}}_{T}={\mathcal{N}}(0,I_{d})$ be the reference Gaussian normal distribution. In diffusion models [ho2020denoising, song2020score], the data distribution is gradually corrupted with Gaussian noise with the perturbation kernel $p({\bm{x}}_{t}|{\bm{x}}_{0})={\mathcal{N}}({\bm{x}}_{0};0,t^{2}I_{d})$, until it approximates a standard normal distribution. Diffusion models learn the denoising generative reverse process, i.e. creating data from noise. This process is governed by the score function $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t})$ [hyvarinen2005estimation], which is estimated through denoising score matching ${\bm{s}}_{\theta}({\bm{x}}_{t})\approx\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t})$ [hyvarinen2005estimation]. Text conditional diffusion models [rombach2022high, deepfloyd_if] are trained with image-text pairs with random dropping so that it is possible to both sample from $p({\bm{x}})$ by using ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t})$ or from $p({\bm{x}}|{\bm{c}})$ by using ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},{\bm{c}})$, where ${\bm{c}}$ is the text embedding. Further, to emphasize the conditioning signal, another widely used technique is classifier free guidance (CFG) [ho2021classifierfree], where ${\bm{\epsilon}}_{\theta}^{\gamma}({\bm{x}}_{t},{\bm{c}}):={\bm{\epsilon}}_{\theta}({\bm{x}}_{t})+\gamma({\bm{\epsilon}}_{\theta}({\bm{x}}_{t},{\bm{c}})-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}))$ is used to sample from the sharpened posterior $p({\bm{x}})p({\bm{c}}|{\bm{x}})^{\gamma}$.
As the loss function, we use the epsilon-matching loss function commonly used in DDPMs [ho2020denoising] and employ a fixed variance setting rather than learning the variance. The model parameters are optimized for 50 epochs using the AdamW optimizer with a learning rate of 0.0001. All experiments were conducted using PyTorch 2.1 in Python with CUDA 12.1, running on 8 NVIDIA H100 GPUs (80GB). The effective batch size was set to 64.
@@ 4.4 Diffusion model architecture
We utilize a pixel-level diffusion model for both unconditional and metadata-conditional generation. The decision to employ pixel-level diffusion, rather than latent-space diffusion via a Variational Autoencoder (VAE) [kingma2013auto], stems from the inherent limitations of VAEs. Specifically, VAE-based reconstruction often compromises high-frequency details [elata2024novel], and solving inverse problems in the latent space requires decoding to ensure data consistency. While such methods [chung2024prompttuning, rout2024solving, kim2023regularization] are theoretically feasible, they frequently fail in practice and lack consensus as a reliable solution. Consequently, we opted to perform generation directly in pixel space.
To train the diffusion model on MRI complex values, we separate the real and imaginary parts, concatenating them along the channel dimension to create a 2-channel input. This results in a 2D input of size 2 × 320 × 320, which is used during training. Our network is adapted from the DeepFloyd IF [deepfloyd_if] pixel-level text-to-image diffusion model, with modifications to accommodate our input format by adjusting the number of channels in the layers. Unlike natural image generation, the textual metadata associated with MRI data is significantly smaller. Therefore, we use the CLIP text encoder instead of the T5-XXL encoder employed in DeepFloyd IF, offering a more compact and memory-efficient representation of the metadata. During training, the CLIP text encoder is frozen, and only the parameters of the UNet modules are learnable.
@@ 4.5 Inverse problem solving with diffusion models
where for the multi-coil MRI case [deshmane2012parallel], the forward model is defined in parallel for each measurement coil
@@
to write the score function of the posterior in terms of the prior score function. A vast literature on DIS can be structured as different ways of approximating the intractable time-dependent likelihood $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$ [daras2024survey]. A few recent works [chung2024prompttuning, kim2023regularization, kim2025dreamsampler] tried to incorporate text conditions through Stable Diffusion [rombach2022high]. Among these, P2L [chung2024prompttuning] is prohibitively slow and is highly unstable when used with CFG. TReg [kim2023regularization] is tailored for severe degradations and produces cartoon-like artifacts. Dreamsampler [kim2025dreamsampler] suggests tailored algorithms for specific types of inverse problems. Generally, setting moderately high values of CFG leads to a steep decrease in the stability of all the algorithms. None of the existing methods have shown the stability and performance required to extend the method to the medical imaging setting, where this is crucial.
We opt for a simple, established baseline for our choice of diffusion model-based inverse problem solver. For this, we choose DDS [chung2024decomposed], which is known to be especially robust to large-scale medical imaging inverse problems. Specifically, with a variance preserving form, DDS with DDIM sampling [song2020denoising] can be described as iterating the following steps for $t=T,T-1,\dots,1$ initialized with random Gaussian noise.
$\hat{\bm{x}}_{0|t}$ $=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left({\bm{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\theta}^{\gamma}({\bm{x}}_{t},{\bm{c}})\right)$ (4)
$\hat{\bm{x}}^{\prime}_{0|t}$ $=\operatorname*{arg\,min}_{{\bm{x}}}\frac{\xi}{2}\|{\bm{y}}-A{\bm{x}}\|_{2}^{2}+\frac{1}{2}\|{\bm{x}}-\hat{\bm{x}}_{0|t}\|_{2}^{2}$ (5)
@@
where $\sigma_{t}:=\eta\sqrt{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ with a constant hyperparameter $\eta=0.8$. From a high-level viewpoint, skipping the data consistency optimization step in Eq. 5 would lead to DDIM sampling from the prior distribution $p({\bm{x}})$. In order to sample from the approximate posterior, DDS modulates the posterior mean $\hat{\bm{x}}_{0|t}={\mathbb{E}}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ with data consistency steps to approximate the conditional posterior mean $\hat{\bm{x}}^{\prime}_{0|t}\approx{\mathbb{E}}[{\bm{x}}_{0}|{\bm{x}}_{t},{\bm{y}}]$. Specifically, for the optimization problem in Eq. 5, the key of DDS is to use few-step conjugate gradient steps. In our case, we use 5-step update with $\xi=5.0$.
Our method is tested on two different undersampling patterns - uniform1D as implemented in fastMRI [zbontar2018fastmri], and poisson2D as implemented in sigpy [ong2019sigpy]. For uniform1D, we test two different acceleration factors: ×$\times 4$ with 8% autocalibrating signal (ACS) region in the center, and ×$\times 8$ with 4% ACS region. For poisson2D, we test ×$\times 8$ and ×$\times 15$ acceleration. For quantitative evaluation, we use three standard metrics - PSNR, SSIM, and LPIPS.
[…]
Figure 6: Metadata statistics of fastMRI knee train set
Figure 7: Metadata statistics of fastMRI brain train set
Figure 8: Metadata statistics of SKM-TEA dataset